OpenStudy (anonymous):

46. Find a function y=f(x) whose second derivative is y" = 12x - 2 at each point (x, y) on its graph and y=- x+5 is tangent to the graph at the point corresponding to x = 1 .

OpenStudy (anonymous):

we know \(f'(x)=6x^2-2x+C\) lets see if we can find \(C\)

OpenStudy (anonymous):

if \(y=-x+5\) is tangent to the curve at \(x=1\) we know that \(f'(1)=-1\) since the slope of \(y=-x+5\) is \(-1\) and since \(f'(1)\) is that slope setting \(f'(1)=6-2+C=-1\) we can solve for \(C\)

OpenStudy (anonymous):

i get \(C=-5\), making \(f'(x)=6x^2-2x-5\)

OpenStudy (anonymous):

this makes \(f(x)=2x^3-x^2-5x+C\) and we have another constant to find. but \(f(x)\) must agree with \(y=-x+5\) at \(x=1\) so replace \(x\) by 1 in both \(f(x)\) and \(y\) and solve for \(C\)